OpenStudy (anonymous):

Algebra Mixture: a chemist has 80 ml of a solution containing 20% acid. How many ml must he romove and replace with pure acid to obtain a 40% solution of acid?

OpenStudy (apoorvk):

well. so i am assuming that is 20% V/V. so, 80 ml has 20% acid - what is the volume of acid in those 80 mls then?

OpenStudy (anonymous):

the question doesn't include the volume

OpenStudy (anonymous):

Do you know the right answer?

OpenStudy (anonymous):

No :(

OpenStudy (mertsj):

Let x = the amount of pure acid to be added .20(80-x)+1.00x=.40(80)

OpenStudy (anonymous):

Tysm you really helped me! I get it now :)

OpenStudy (mertsj):

yw. What is the answer?

OpenStudy (anonymous):

x=20

OpenStudy (mertsj):

Good job!!

OpenStudy (anonymous):

thnx xD

OpenStudy (anonymous):

 \[\frac 14 \times 80 = 20 L \]